A neutron in a reactor makes an elastic headon collision with the nucleus of an atom initially at rest.
Assume: The mass of the atomic nucleus is
about 11.1 the mass of the neutron.
What fraction of the neutron’s kinetic energy is transferred to the atomic nucleus?
Don't round answer.
1
Expert's answer
2019-12-04T08:44:46-0500

When the collision is elastic the energy of the system and total momentum is conserved so


"E(n0) = E(n) +E(nuc)""P(n0)=P(n)+P(nuc)""E(n0)=(m(n)v(n0)^2)\/2""E(n)=(m(n)v(n)^2)\/2""E(nuc)=(m(nuc)v(nuc)^2)\/2=(11.1m(n)v(nuc)^2)\/2""P(n0)=m(n)v(n0)""P(n)=m(n)v(n)""P(nuc)=m(nuc)v(nuc)=11.1m(n)v(nuc)""P(n)=P(n0)-P(nuc)=m(n)v(n0)-11.1m(n)v(nuc)=m(n)v(n)""v(n) =(v(n0)-11.1v(nuc))"

from the first equation


"(m(n)v(n0)^2)\/2=(m(n)(v(n0)-11.1v(nuc))^2)\/2+(11.1m(n)v(nuc)^2)\/2""v(n0)^2=(v(n0)-11.1v(nuc))^2+11.1v(nuc)^2""v(n0)^2=v(n0)^2-22.2v(n0)v(nuc)+123.21v(nuc)^2+11.1v(nuc)^2""22.2v(n0)v(nuc)=134.31v(nuc)^2""v(nuc)=(22.2\/134.31)v(n0)=(1\/6.05)v(n0)"

the fraction of the neutron energy transmitted to the nucleus can be calculated as


"f=E(nuc)\/E(n0)=(11.1m(n)(((1\/6.05v)(n0))^2))\/(m(n)v(n0)^2)=0.303"